# LINKAGE: Listwise Ranking among Varied-Quality References for Non-Factoid QA Evaluation via LLMs

### v1

@@ Abstract
Non-Factoid (NF) Question Answering (QA) is challenging to evaluate due to diverse potential answers and no objective criterion. The commonly used automatic evaluation metrics like ROUGE or BERTScore cannot accurately measure semantic similarities or answers from different perspectives. Recently, Large Language Models (LLMs) have been resorted to for NFQA evaluation due to their compelling performance on various NLP tasks. Common approaches include pointwise scoring of each candidate answer and pairwise comparisons between answers. Inspired by the evolution from pointwise to pairwise to listwise in learning-to-rank methods, we propose a novel listwise NFQA evaluation approach, that utilizes LLMs to rank candidate answers in a list of reference answers sorted by descending quality. Moreover, for NF questions that do not have multi-grade or any golden answers, we leverage LLMs to generate the reference answer list of various quality to facilitate the listwise evaluation. Extensive experimental results on three NFQA datasets, i.e., ANTIQUE, the TREC-DL-NF, and WebGLM show that our method has significantly higher correlations with human annotations compared to automatic scores and common pointwise and pairwise approaches.
@@
Sihui Yang1,2 Keping Bi1,2 Wanqing Cui1,2 Jiafeng Guo1,2 Xueqi Cheng1,2 1CAS Key Laboratory of Network Data Science and Technology, Institute of Computing Technology, Chinese Academy of Sciences; 2School of Computer Science and Technology, University of Chinese Academy of Sciences; {yangsihui22s, cuiwanqing18z, bikeping, guojiafeng, cxq}@ict.ac.cn,
@@ 5.1 Overall Results
Few-shot in-context Learning can enhance the performance of LINKAGE. Comparing with results under few-shot and zero-shot, providing LLMs with a few examples can help demonstrate the evaluation task more clearly. Compared to Mistral, the enhancement of few-shot ICL on ChatGPT is less. We think that it is because ChatGPT has a much better understanding of instructions so the few-shot example does not help it much. More experiment details about few-shot learning are in the Appendix B.1.
@@ 5.3 Study on the Reference List Randomness
Our experiments involve random sampling of the ground truth set, so we evaluate the results under 3 randomizations to analyze the impact of randomness on performance. Details about experiments are in Appendix B.2. From Figure 3, we can observe that for all LLMs on all datasets, the standard deviations of the experiments are always small. This indicates that the randomness of the selection of reference answers has little impact on the evaluation results, which proves that the improvement brought by our method is significant.
@@ A.2 Instruction for Generating Reference List
Appendix B Experiment Details
B.1 N-shot LINKAGE performance
We conduct several sets of few-shot experiments of LINKAGE on TREC-DL-NF and ANTIQUE using Mistral. Results are in Table 7 and Table 8. We find the number of samples cannot be too large. When the number exceeds a certain value, the performance will deteriorate. This is because the shot number increasing leads to a significant increase in the input length, which will make the LLMs difficult to understand.
n-shot Kendal Spearman Pearson
$n$=0 0.3287 0.3539 0.3401
$n$=1 0.4246 0.4656 0.4724
$n$=3 0.4312 0.4752 0.4958
$n$=5 0.4339 0.4725 0.4958
n-shot Kendal Spearman Pearson
$n$=0 0.2900 0.3101 0.3172
$n$=1 0.3850 0.4183 0.4256
$n$=3 0.3696 0.4012 0.4122
$n$=5 0.3654 0.3934 0.4041
B.2 Experiments on Randomness of Reference List
Three independent experiments on randomly selecting $\mathcal{R}$ on TREC-DL-NF and ANTIQUE using ChatGPT and Mistral are shown in Table 9 and Table 10.
Model Method Spearman 1 Spearman 2 Spearman 3 Average Std
Mistral PointwiseR≠∅ 0.3382 0.3463 0.3567 0.3471 0.0093
Pairwise 0.2967 0.2783 0.2912 0.2887 0.0094
LINKAGEfew_shot 0.4725 0.4579 0.4520 0.4608 0.0105
ChatGPT PointwiseR≠∅ 0.2777 0.4288 0.4526 0.3864 0.0948
Pairwise 0.3692 0.3687 0.3544 0.3641 0.0083
LINKAGEfew_shot 0.4094 0.3854 0.4325 0.4091 0.0235
Table 10: Average Spearman coefficient and standard deviation of randomly select $\mathcal{R}$ in three dependent experiments on ANTIQUE using Mistral and ChatGPT.

### v2

@@ Abstract
Non-Factoid (NF) Question Answering (QA) is challenging to evaluate due to diverse potential answers and no objective criterion. The commonly used automatic evaluation metrics like ROUGE or BERTScore cannot accurately measure semantic similarities or answers from different perspectives. Recently, Large Language Models (LLMs) have been resorted to for NFQA evaluation due to their compelling performance on various NLP tasks. Common approaches include pointwise scoring of each candidate answer and pairwise comparisons between answers. Inspired by the evolution from pointwise to pairwise to listwise in learning-to-rank methods, we propose a novel listwise NFQA evaluation approach, that utilizes LLMs to rank candidate answers in a list of reference answers sorted by descending quality. Moreover, for NF questions that do not have multi-grade or any golden answers, we leverage LLMs to generate the reference answer list of various quality to facilitate the listwise evaluation. Extensive experimental results on three NFQA datasets, i.e., ANTIQUE, the TREC-DL-NF, and WebGLM show that our method has significantly higher correlations with human annotations compared to automatic scores and common pointwise and pairwise approaches. Our code and dataset can be found at https://github.com/babyyang525/LINKAGE-Listwise-NFQA-Evaluation.
@@
Sihui Yang1,2 Keping Bi1,2thanks: Corresponding author. Wanqing Cui1,2 Jiafeng Guo1,2 Xueqi Cheng1,2 1CAS Key Laboratory of Network Data Science and Technology, Institute of Computing Technology, Chinese Academy of Sciences 2University of Chinese Academy of Sciences, Beijing, China {yangsihui22s, bikeping,cuiwanqing18z, guojiafeng, cxq}@ict.ac.cn
@@ 5.1 Overall Results
Few-shot in-context Learning can enhance the performance of LINKAGE. Comparing with results under few-shot and zero-shot, providing LLMs with a few examples can help demonstrate the evaluation task more clearly. Compared to Mistral, the enhancement of few-shot ICL on ChatGPT is less. We think that it is because ChatGPT has a much better understanding of instructions so the few-shot example does not help it much.
n-shot Kendal Spearman Pearson
$n$=0 0.3287 0.3539 0.3401
$n$=1 0.4246 0.4656 0.4724
$n$=3 0.4312 0.4752 0.4958
$n$=5 0.4339 0.4725 0.4958
n-shot Kendal Spearman Pearson
$n$=0 0.2900 0.3101 0.3172
$n$=1 0.3850 0.4183 0.4256
$n$=3 0.3696 0.4012 0.4122
$n$=5 0.3654 0.3934 0.4041
However, the number of samples cannot be too large. We conduct several sets of few-shot experiments on TREC-DL-NF and ANTIQUE using Mistral. As the results are shown in Table 7 and Table 8. When the number exceeds a certain value, the performance will deteriorate. This is because the shot number increasing leads to a significant increase in the input length, which will make the LLMs difficult to understand.
Model Method Spearman 1 Spearman 2 Spearman 3 Average Std
Mistral PointwiseR≠∅ 0.3382 0.3463 0.3567 0.3471 0.0093
Pairwise 0.2967 0.2783 0.2912 0.2887 0.0094
LINKAGEfew_shot 0.4725 0.4579 0.4520 0.4608 0.0105
ChatGPT PointwiseR≠∅ 0.2777 0.4288 0.4526 0.3864 0.0948
Pairwise 0.3692 0.3687 0.3544 0.3641 0.0083
LINKAGEfew_shot 0.4094 0.3854 0.4325 0.4091 0.0235
@@ 5.3 Study on the Reference List Randomness
Our experiments involve random sampling of the ground truth set, so we evaluate the results under 3 randomizations to analyze the impact of randomness on performance. Results on TREC-DL-NF are in Table 9 and results on ANTIQUE can be found in Appendix B. We can observe that for all LLMs on all datasets, the standard deviations of the experiments are always small. This indicates that the randomness of the selection of reference answers has little impact on the evaluation results, which proves that the improvement brought by our method is significant.
@@ 7 Conclusion
Acknowledgement
This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No. 62302486, the Innovation Project of ICT CAS under Grants No. E361140, the CAS Special Research Assistant Funding Project, the Lenovo-CAS Joint Lab Youth Scientist Project, the project under Grants No. JCKY2022130C039, and the Strategic Priority Research Program of the CAS under Grants No. XDB0680102.
@@ A.2 Instruction for Generating Reference List
Appendix B Experiments on Randomness of Reference List
Three independent experiments on randomly selecting $\mathcal{R}$ on ANTIQUE using ChatGPT and Mistral are shown in Table 10.
Table 10: Results of randomly select $\mathcal{R}$ in three dependent experiments on ANTIQUE.